OpenStudy (anonymous):

Write the equation that is parallel to Y=1/2x + 3and passes through (-6,5)

OpenStudy (anonymous):

You get the slope from the first line, then use the formula for getting the line while knowing the point and slope.

OpenStudy (radar):

And knowing that a parallel line will have the same slope.

OpenStudy (radar):

What is the slope of the given line?

OpenStudy (radar):

Do you know where to look?

OpenStudy (anonymous):

ummm

OpenStudy (radar):

ummm is not the answer. (From Judge Judy)

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

1/2 is the slope

OpenStudy (radar):

The given line y=(1/2)x + 3 The equation is already in slope intercept form, so just look at the coefficient of the x term, that is the slope.

OpenStudy (anonymous):

so what do i do next?

OpenStudy (radar):

So the new line to be parallel will also have the slope 1/2. Now use the fact that slope is also equal to delta y/ delta x or\[\Delta y \over \Delta x\]

OpenStudy (radar):

(y-5)/(x+6) =1/2 that is also delta y/delta x

OpenStudy (radar):